Operation Summary for September 23, 2011 - September 26, 2011 |
The Integrated Luminosity for the period from 9/19/11 to 9/26/11 was 58.32 pb-1 (inverse picobarns). NuMI reported receiving 2.1E18 protons on target during this same period. |
Friday September 23 The day shift began with Operations monitoring the stash (381.7E10), with Pbar stacking (21.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target cooling water tests, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting on the next Tevatron (TEV) store. During the day shift, Operations put the TEV into HEP shot setup at 8:07 AM. The Neutron Therapy Facility began tuning their beamline at 8:09 AM in preparation for treating one patient later in the morning. Operations established store 9147 at 9:06 AM with an average initial luminosity (AIL) of 346.4E30. Operations established beam to the Meson FTBF experiment T-1017 at 9:52 AM. Operators reported that during this shift they had rebooted MECAR and that MI personnel had conducted coalescing studies. During the evening shift, at 7:05 PM, operators reported that the Tevatron Electron Lens (TEL2) vacuum pump tripped off and would not reset; operators informed people. Operations disabled beam to T-1017 at 7:12 PM; this ended T-1017’s run. Saturday September 24 The midnight shift began with Operations monitoring store 9147 (75.03E30) and the stash (367E10), with Pbar stacking (6.8E10), and with MiniBooNE taking beam. During the midnight shift, at 1 AM, a Recycler kicker (R:KPS3A) tripped off due to a low Fluorinert level; operators called in people to fill the kicker after the next store was established. Operations terminated store 9147 at 1:52 AM, put the TEV into shot setup, and established store 9148 at 3 AM with an AIL of 335.99E30. Operators prepared MI for access at 3:18 PM. Operators reported at 4:21 AM that they had been having problems with a Booster RF station (BRF1) for the last hour; tuning seemed to help the station put out more RF. MI came out of access: Pbar resumed stacking and MiniBooNE resumed taking beam by 5:05 AM. During the day shift, Operations established beam to NuMI at 9:50 AM and the beam permit immediately tripped off due to a high current on a quadrupole magnet power supply (E:QF119); it reset. Operators reported that BRF1 was again causing problems; operators adjusted its anode voltage. Operations terminated store 9148 at 3:05 PM. During the evening shift, at 4:04 PM, the MI South Anode Power Supply tripped off; it reset. At 4:22 PM, MIRF1 tripped off and tripped off the MI beam permit for the first of two times in one hour and 45 minutes; it reset. Operators put the TEV through a wet squeeze at 6:28 PM. Operations put the TEV into shot setup at 6:48 PM and established store 9151 at 8 PM with an AIL of 313.98E30. At 8:36 PM, the I- Source vacuum valve closed due to bad vacuum in tank 1; it reset and opened. Sunday September 25 The midnight shift began with Operations monitoring store 9151 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 7:19 AM, operators noticed the H- Source high voltage (the backup source) had tripped off; operators reset it. During the day shift, operators reported at 8:53 AM that the TEV sector C1 electronics room air conditioning (AC) unit was having trouble; FESS personnel installed a portable AC unit. At 10:29 AM, a Linac Klystron RF station (KRF2) tripped off for the first of two times in 13 minutes; it reset. Operations terminated store 9151 at 10:47 AM, put the TEV into shot setup, and established store 9152 at 12:08 PM with an AIL of 348.91E30. During the evening shift, at 7:26 PM, a Booster RF vacuum pump (B:IPRF19) tripped off and brought down BRF17 and 18; everything reset. The NuMI beam permit tripped off at 8:22 PM when the output of a power supply (E:V118) dropped low; everything reset. Operators tuned the Linac and Booster for better efficiency at 11 PM. Monday September 26 The midnight shift began with Operations monitoring store 9152 and the stash, with Pbar stacking, and with the NuMI and MiniBooNE experiments taking beam. During the midnight shift, operators performed a MI coalescing check; everything looked good. Operations terminated store 9152 at 3:45 AM, put the TEV into shot setup, and established store 9153 at 5:04 AM with an AIL of 343.08E30. The NuMI beam permit tripped off due to E:V118 drifting low; everything reset. Operators reported short downtimes due to problems with LRF1 and 5, KRF5, MIRF9, and to bad AP1transfer line beam positions. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster aperture scans. |
|
More Information Comments and Suggestions |
| last modified 09/26/2011 email Fermilab |
| Security, Privacy, Legal | |